I don't know how to figure this out On thursday marie worked for 6 hours and handled 20 calls. On friday marie worked 8 hours and handled 30 calls. If she worked 4 hours how many calls did she take? Please help I think the m=5 but i'm not sure what to do next.
Answer by stanbon(57347)

On thursday marie worked for 6 hours and handled 20 calls. On friday marie worked 8 hours and handled 30 calls. If she worked 4 hours how many calls did she take?
---------
You have two points: (6,20) and (8,30)
slope = (30-20)/(8-6) = 5
Equation form is: calls = (slope)(hours)+b
Substitute to find "b", as follows:
20 = 5*6 + b
b = -10
EQUATION:
calls = 5(hours)-10
------------
Let # of hours be 4
Then # of calls = 5*4-10 = 10
--------------
Marie took 10 calls
==============
Cheers,
Stan H.